# Fast, accurate training and sampling of Restricted Boltzmann Machines

### v1

@@ 4 The low-rank RBM pretrained
In Ref. [41], it was shown that it is possible to train exactly (i.e. by direct numerical integration instead of MCMC sampling) an RBM containing a reduced number of modes in the weight matrix $\bm{W}$ by exploiting a mapping between the RBM and a Restricted Coulomb Machine and solving a convex optimization problem, see the SI. In other words, it is possible to train a RBM with a coupling matrix of this simplified form
@@
and where the right singular vectors $\{\bm{u}_{\alpha}\}_{\alpha=1}^{d}$ correspond exactly to the first $d$ principal directions of the data set. Under this assumption, it is possible to write $p(\bm{v})$ only as a function of $d$ order parameters given by the magnetizations along each of the $u_{\alpha}$ components, $m_{\alpha}(\bm{v})=\bm{u}_{\alpha}\cdot\bm{v}/\sqrt{{N_{\mathrm{v}}}}$, and in particular,
@@
$H(\bm{v})=-\sum_{a}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\bar{u}_{a}\sum_{\alpha=1}^{d}w_{\alpha}m_{\alpha}+\eta_{a}\right)=\mathcal{H}(\bm{m}(\bm{v})),$ (4)
@@
where $\bm{m}=\left(m_{1},\ldots,m_{\alpha}\right)$. As proposed in [41], the optimal parameters of such a model can basically be determined by solving a regression problem. We describe this method in details in the SI. This means that once the model is trained, we obtain a probability $p(\bm{m})$ defined on a much lower dimension than the original $p(\bm{v})$. Such a probability can be straightforwardly sampled using inverse transform sampling. Since this method requires a discretization of the $\bm{m}$-space both for training and generation, we cannot consider intrinsic space dimension $d>4$ dimensions in practice. These low-rank RBMs are then trained to reproduce the statistics of the dataset projected in its first $d$ principal components. Despite their simplicity, the low-rank models are already able to generate an approximate version of the dataset, as shown in Fig. 1–D-F for the 4 datasets previously presented.
@@
In the initial stage of the standard learning process, the model encodes the strongest PCA components of the data through multiple critical transitions [26, 27, 28]. Pre-training with the low-rank construction allows us to bypass these transitions and avoid out-of-equilibrium effects caused by critical slowing down associated to these transitions. Once the main directions are incorporated, training can efficiently continue with standard algorithms like PCD, as the mixing times of pre-trained machines tend to be much shorter. In particular, in the PCD-100 training with MNIST01, relaxation times for the visible variables’ time correlation reach $5\cdot 10^{5}$ MCMC steps at the first three transitions, coinciding with the growth of singular values in the model weight matrix $W$. In contrast, the pre-trained machine has a much shorter relaxation time of $\sim 10^{3}$, allowing us to safely restart the PCD process from a set of equilibrium samples generated by static sampling of the low-rank RBM.
Figure 2: We compare the equilibrium samples generated by RBMs trained on the Mickey, MNIST01, and HGD datasets using three different training schemes: Jarzynski (JarRBM), PCD, and PCD initialized on low-rank RBMs (used to generate the samples in Fig. 1–D-F). To assess the fitting of the modes, we show a density plot of the projections of the data in the first two principal directions of each dataset. We compare these results with the density plot of the original datasets in the first column.
@@
Overcoming these transitions has dramatic implications for the quality of the models we can train and how accurately they reproduce the statistics of the data. In Fig. 2, we show for 3 datasets the equilibrium samples drawn from 3 RBMs trained with identical number of samples, minibatch size, $k=100$ Gibbs steps, and learning rate $\gamma=0.01$, but different training strategies. In particular, we consider 2 RBMs trained from scratch with the standard PCD [33] and the recently proposed Jarzynski reweighing method [23] (see SI for our specific implementation in the RBM), and a final machine trained with PCD and pre-trained with a low-rank RBM. In all cases, the quality of the generated samples is significantly better when pre-training is used. For the Mickey dataset, neither JarRBM nor normal PCD are able to generate convincing data. For the MNIST01 dataset, all 3 methods are able to generate convincing data, but only Pretrain+PCD is able to correctly balance all modes, as can be seen in Fig. 3, where we compare the histograms of the generated data projected onto the first 3 PCA directions with those of the dataset and a random selection of the generated samples. We see that the pre-training+PCD training perfectly balances the different modes (here we show the first 3 directions, but it goes much further), unlike the other 2 methods, and also generates more diverse images. We can also compare the log-likelihood of all 3 models and find that the pre-trained RBM achieves higher values. At this point, it is important to emphasize that in order to properly quantify the increase in log-likelihood, we need to use the PTT algorithm (see section 5) to correctly thermalize in these well-trained machines. For comparison, we show our PTT measure in dark and solid lines, while the standard AIS [42] estimate is shown in light dashed lines.
Figure 3: We compare the samples generated by the 3 RBMs (JarRBM, PCD, pretrain+PCD) trained with MNIST01 data. In A, we show the histograms of the generated data projected on the first, second and third principal directions with those of the dataset. We see that only the pretrain+PCD correctly balances the different modes. In B we show 10 images generated by each machine. In C, we compare the log-likelihood of each model’s dataset as a function of training time. The dark and full curves were obtained using the PTT algorithm discussed in section  5, and the lighter and dashed curves using the AIS method [42].
Already from the scatter plots we see that the pre-training has a dramatic effect in obtaining models where all modes are properly balanced, but also has important effects in the maximum test-likelihood we can achieve. In all cases, these equilibrium samples are drawn using the trajectory PT algorithm that will be explained in the next section, and the log-likelihood obtained using the equilibrium configurations obtained at different epochs as a result of the trajectory PT flow.
5 Standard Gibbs sampling vs. Parallel Trajectory Tempering (PTT)
@@
One major challenge with structured datasets is quantifying the model’s quality, since sampling the equilibrium measure of a well-trained model is often too time-consuming. This affects the reliability of generated samples and indirect measures as log-likelihood’s estimation through Annealing Importance Sampling (AIS) [42], making them inaccurate and meaningless.
Figure 4: Comparison between PTT and classical Gibbs sampling for the MNIST01 dataset (A and B, respectively) and the human genome dataset (C and D, respectively). In A and C, we show the trajectory of two independent chains (red and orange) projected onto the PCA along the sampling process of the pretraining+PCD model for $10^{4}$ MCMC steps. The black contour represents the density profile of the dataset and the position of the chains is plotted every 10 steps. In B and D we show the average number of jumps from one cluster to another as a function of the MCMC steps performed. The average is calculated over a population of 100 chains. In D, we show the average jump time between clusters along the first (solid line) and second (dashed line) principal components of the data.
@@
To illustrate this problem, let us consider the MNIST01 and the HGD datasets. MNIST01 dataset is bimodal and the HGD highly multimodal as shown in their PCA in Figs. 1–A and C. Let us consider that we want to sample the equilibrium measure of the RBMs trained using low-rank RBM pretraining. In order to draw new samples from these models, one would typically run MCMC processes from random initialization and iterate them until convergence. The mixing time is controlled by the jumping time between clusters. To accurately estimate the relative weight between modes, the MCMC processes must be ergodic, requiring many back-and-forth jumps. However, as shown in Figs. 4–A and C for the MNIST01 and HGD datasets, Gibbs sampling dynamics are extremely slow, rarely producing jumps even after $10^{4}$ MCMC steps. The yellow curves in Figs. 4–B and D show the mean number of jumps over 100 independent chains as a function of MCMC steps, indicating that a proper equilibrium generation would require at least $10^{6}-10^{7}$ MCMC steps.
One effective way to accelerate the dynamics is to exploit the training trajectory, where the model progressively specializes through second-order phase transitions. To achieve this, we save RBMs trained at various epochs and propose swaps between configurations of similarly trained models. We call this the Parallel Trajectory Tempering (PTT) algorithm. Unlike the standard Parallel Tempering (PT) algorithm, which attempts swaps configurations between different temperatures, the PTT swaps between model parameters with different degrees of specialization. This approach is more natural for this problem because it involves crossing only second-order transitions, unlike the first-order transitions occurring in temperature annealing. And in fact, we show in Figs. 4–A and C, that this approach allows us to sharply accelerate the dynamics, as opposed to the standard PT algorithm (studied in detail for the MNIST dataset in [40]).
In the PTT algorithm, the configurations $\bm{x}=(\bm{v},\bm{h})$ of neighboring machines indexed by $t$ and $t-1$ are interchanged with the probability
$p_{\mathrm{acc}}(\bm{x}^{t}\leftrightarrow\bm{x}^{t-1})=\min\left(1,\exp\left(\Delta\mathcal{H}^{t}(\bm{x}^{t})-\Delta\mathcal{H}^{t}(\bm{x}^{t-1})\right)\right).$
This move satisfies detailed balance with our target equilibrium distribution $p(\bm{x})=\exp(-\mathcal{H}(\bm{x}))/Z$, ensuring that the moves lead to the same equilibrium measure. As “nonspecialized" models mix very quickly, either because the distribution is essentially Gaussian at the initialisation of a standard training, or because the low-rank RBM can be sampled with a static Monte Carlo process (yielding independent configurations each time), the trajectory flow significantly accelerates convergence to equilibrium. The time interval between successive machines is selected in such a way that the probability of accepting interchanges between neighboring machines remains around 0.3. Pre-trained machines require a significant fewer number of models to be effective, because most selected models are positioned at the most prominent phase transitions. We give the number of machines used for each sampling process in the SI. We also provide there a specific and detailed description of the algorithm used.
In the red curves in Fig. (4)–B and D, we show the number of jumps between clusters as a function of the number of elementary MCMC steps, which in the PTT scheme refer to 1 Gibbs sampling step + one swap proposal. For the DNA dataset, we have two measures corresponding to jumps along the two principal component directions. We observe at $10^{4}$ MCMC steps an increase of the number of jumps by a factor of 80 for MNIST01 and by a factor of 1350 for the HGD in this machine, although we achieve higher factors in other machines, as we show in the SI. The sampling of RBMs training on the MNIST01 dataset was the subject of the study of the “stacked tempering" algorithm in  [40]. If we compare the numbers with their work, we see that we achieve a 3-4 times higher speedup factor, where our model has the advantage that it does not need additional training, but simply uses the stored machines correctly.
Another desirable advantage of our PTT algorithm is that we can easily use it to compute an improved estimate of the AIS log-likelihood, except that in our case we consider the training trajectory instead of a cooling process and use the equilibrium samples obtained for each of the models to compute the model averages. In Figs. 3–C  5–A we compare the log-likelihood estimates obtained with our method (AIS-PTT) in full and dark lines and in light and dashed lines the AIS estimate (AIS). We see that both measures coincide for most parts of the training and that they split when the sampling becomes too long to thermalize along the temperature annealing curve in AIS. This effect is particularly evident for the JarRBM run in  5–A, where AIS takes a long time to recognize that the model suffers from a strong mode-collapse effect.
6 Overfitting and privacy loss as quality indicators
In this section, we examine the quality of the samples generated, regarding overfitting and privacy criteria which have been defined for genomic data in particular. We look at this on the models trained with PCD with and without pre-training. We do not include the Jarzysnki method here, as this method fails to obtain a reliable model as clearly shown in the evolution of the Log-likelihood in Fig. 5. We focus on the human genome dataset, as shown in Fig. 1–C, to evaluate the ability of various state-of-the-art generative models to generate realistic fake genomes while minimizing privacy concerns (i.e., reducing overfitting). Recent studies [17, 18] have thoroughly investigated this for a variety of generative models. Both studies concluded that the RBM was the most accurate method for generating high-quality and private synthetic genomes. The comparison between models relies primarily on the Nearest Neighbor Adversarial Accuracy ($AA_{\rm TS}$) and privacy loss indicators, introduced in Ref. [43], which quantify the similarity and the level of "privacy" of the data generated by a model w.r.t. the training set. We have $AA_{\rm TS}=\frac{1}{2}\bigl{(}AA_{\rm True}+AA_{\rm Synth}\bigr{)}$ where $AA_{\rm True}$ [resp. $AA_{\rm Synth}$] are two quantities in $[0,1]$ obtained by merging two sets of real and synthetic data of equal size $N_{s}$ and measuring respectively the frequency that a real [rep. synthetic] has a synthetic [resp. real] as nearest neighbor. If the generated samples are statistically indistinguishable from real samples, both frequencies $AA_{\rm True}$ and $AA_{\rm Synth}$ should converge to 0.5 at large ${N_{\mathrm{s}}}$. $AA_{\rm TS}$ can be evaluated both with train or test samples and the privacy loss indicator is defined as ${\rm Privacy\ loss}=AA_{\rm TS}^{\rm test}-AA_{\rm TS}^{\rm train}$ and is expected to be strictly positive. Fig. 5 shows the comparison of $AA_{\rm TS}$ and privacy loss values obtained with our two models, demonstrating that the pre-trained RBM clearly outperforms the other model, and even achieves better results ($AA_{\rm TS}$ values much closer to 0.5) than those discussed in [17, 18].
Figure 5: We compare the quality of the RBMs trained with the human genome data (HGD). In A, we show the log-likelihood as a function of the training epochs for the 3 training procedures. Solid lines correspond to AIS-PTT and dashed lines to AIS. The JarRBM falls down because the training breaks eventually. In B and C we compare privacy and overfitting based on the $AA_{\rm TS}$ indicator.

### v2

@@ 4 The low-rank RBM pretrained
In Decelle & Furtlehner (2021a), it was shown that it is possible to train exactly an RBM containing a reduced number of modes in the weight matrix $\bm{W}$ by exploiting a mapping between the RBM and a Restricted Coulomb Machine and solving a convex optimization problem, see the SI A.2. In other words, it is possible to train an RBM whose coupling matrix has this simplified form
@@
where $\{w_{\alpha}\}$ are the singular values of the coupling matrix and the right singular vectors $\{\bm{u}_{\alpha}\}_{\alpha=1}^{d}$ correspond exactly to the first $d$ principal directions of the data set. Under this assumption, $p(\bm{v})$ a function of only $d$ order parameters, given by the magnetizations along each of the $\boldsymbol{u}_{\alpha}$ components, i.e. $m_{\alpha}(\bm{v})=\bm{u}_{\alpha}\cdot\bm{v}/\sqrt{{N_{\mathrm{v}}}}$. In particular,
@@
$H(\bm{v})=-\sum_{\alpha=0}^{d}\theta_{\alpha}m_{\alpha}-\sum_{a}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\bar{u}_{a}\sum_{\alpha=1}^{d}w_{\alpha}m_{\alpha}+\eta_{a}\right)=\mathcal{H}(\bm{m}(\bm{v})),$ (4)
where $\bm{m}=\left(m_{1},\ldots,m_{\alpha}\right)$ and $\theta_{\alpha}$ is the projection of the visible bias on the extended $\bm{u}$ matrix. In order to initialize the visible bias as suggested in Montavon & Müller (2012), a direction $\bm{u}_{0}$ is used for the magnetization $m_{0}$ that is only present in the bias term. This is an improvement of this work with respect to Decelle & Furtlehner (2021a), as it allows us to gain almost for free one additional direction.
@@
The optimal parameters can be determined by solving a convex optimization problem, detailed in SI A.2. Once the model is trained, we obtain a probability distribution $p(\bm{m})$ in a much lower-dimensional space than the original $p(\bm{v})$, which can be easily sampled globally using the inverse transform sampling method. However, since this approach requires discretizing the $\bm{m}$-space for both training and generation, it becomes impractical for intrinsic space dimensions $d>4$ due to memory constraints. These low-rank RBMs are trained to replicate the statistics of the dataset projected onto its first $d$ principal components. Despite their simplicity, the low-rank models can still generate an approximate version of the dataset, as shown in Supplemental Figure 15 for the five datasets previously presented.
@@
In the initial stage of the standard RBM learning process, the model encodes the strongest PCA components of the data through multiple critical transitions (Decelle et al., 2017; 2018; Bachtis et al., 2024). Pre-training with the low-rank construction allows us to bypass these transitions and avoid out-of-equilibrium effects caused by critical slowing down associated to these transitions, which becomes prohibitive in the case of very clustered datasets. Once the main directions are incorporated, training can efficiently continue with standard algorithms like PCD, as the mixing times of pre-trained machines tend to be much shorter that at the transitions.
Figure 2: We compare the equilibrium samples generated by RBMs trained on the MNIST01, Mickey, HGD, Ising2D and CelebA datasets using two different training schemes: PCD, and PCD initialized on low-rank RBMs. To assess the fitting of the modes, we show a scatter plot of the projections of the samples in the first two principal directions of each dataset, together with the histograms along each direction. Black dots and histograms refer to the original data distribution, while the red color is used to represent the generated data. In D we show the evolution of the log-likelihood in time for the two training methods computed with the online AIS trajectory. In continuous lines we show the train and in dashed, the test. Results are obtained by averaging with respect to 10 trainings initialized at the same low-rank RBM, and the standard deviation between all is shown in a shade around the line, which is hardly seen due to the small variation from run to run.
@@
Overcoming these transitions has dramatic implications for the quality of the models we can train and how accurately they reproduce the statistics of the data, especially for short training times. In Fig. 2, we show for 5 datasets the equilibrium samples drawn from 2 RBMs trained with identical number of samples, minibatch size, $k=100$ Gibbs steps, and learning rate $\gamma=0.01$, but different training strategies. In particular, we consider an RBMs trained from scratch with the standard PCD (Tieleman, 2008), and a final machine trained with PCD and pre-trained with a low-rank RBM. We also considered machines trained with the recently proposed Jarzynski reweighing method (Carbone et al., 2023) (see SI E for our specific implementation in the RBM), but we do not include the analysis here because in most of the datasets the training method eventually becomes very unstable reaching very poor results. In all datasets, the quality of the generated samples and the test log-likelihood are significantly better when pre-training is used. We see that the pre-training+PCD training perfectly balances the different modes (here we show the projections on the first 2 PCA directions, but the agreement extends further to several other principal directions). We can also compare the log-likelihood of all 5 datasets and find that the pre-trained RBM achieves higher values. In all cases, the log-likelihood is measured using the annealed importance sampling trajectory (Tr-AIS) method, and the samples are drawn using the trajectory PT (PTT) algorithm. Both algorithms will be described in the next section 5.
Already from the scatter plots, we see that the pre-training is key in obtaining models where all modes are properly balanced in clustered datasets, but it’s also beneficial in the maximum test log-likelihood we can achieve. We observe that the pretraining is not particularly useful when the dataset is not much clustered, as it is the case of CelebA and the entire MNIST dataset as shown in Fig. 16 of the SI I. This result can be easily understood as the encoding of the first modes of the data do not involve splitting the distribution into two isolated modes, something that happens in the case of clustered data.
5 Trajectory annealing
The evolution of parameters during training can be seen as a smoother annealing process compared to the standard temperature annealing used in algorithms like annealing importance sampling (AIS) for log-likelihood estimation or the PT algorithm for sampling. This is because it avoids the first-order transitions that typically occur with temperature in clustered datasets. However, these methods can still be applied to trajectory annealing by using the parameters from different stages of training, rather than scaling the model parameters by a common factor $\beta_{i}$, as is done in temperature-based protocols.
5.1 Trajectory Annealing importance sampling
The standard method for computing the log-likelihood (LL) in the literature is Annealed Importance Sampling (AIS) (Neal, 2001). For a detailed review of its application in RBMs, see Krause et al. (2020). In this approach, the partition function of a model with parameters $\bm{\Theta}=\{\bm{w},\bm{\theta},\bm{\eta}\}$ is estimated by gradually increasing the inverse temperature $\beta$ from 0 to 1 over $N_{\beta}$ steps. At $\beta=0$ (where $p_{0}=p_{\mathrm{ref}}$), the partition function $Z_{0}$ is tractable, enabling the estimation of $Z_{i}$ at neighboring $\beta_{i}$ values through path sampling methods. The reference configuration $p_{\mathrm{ref}}$ is commonly taken uniform, but improvements have been reported in RBM when fixing it to the independent site distribution that describes the empirical center of the dataset. For a detailed description of the algorithm, see SI D.
AIS is quite costly because it requires a large number of annealing steps $N_{\beta}$ for reliable estimations, making it impractical to compute the log-likelihood online during training. As a result, it is often calculated offline using a set of models at different temperatures. However, since the training process itself follows an annealing trajectory, it makes more sense to use the evolution of parameters during training as annealing steps rather than relying on temperature. Additionally, at the start of training, parameters are set to zero, making $Z_{0}$ tractable, which also holds when initializing with a low-rank RBM. Since updating the partition function only involves calculating the energy difference between the new and old parameters, this can be efficiently computed online during each parameter update, allowing for very small integration steps. We refer to this method as online Tr-AIS. Alternatively, similar to standard AIS, multiple models can be saved during training for an offline Tr-AIS estimation.
To evaluate the reliability of our new approach, we trained several RBMs using datasets with a maximum of 20 hidden nodes, allowing for the exact computation of the partition function through direct enumeration. We then calculated the difference between the true partition function and the estimations obtained from various methods. The results for the HGD dataset are presented in Fig. 3, while similar qualitative results for the other datasets can be found in the SI, see Figs. 11, 12, 13 and 14. In Fig. 3, we display the estimation errors in the LL using three different methods: standard temperature Annealed Importance Sampling (AIS) (panel A), AIS with a reference distribution (panel B), and Tr-AIS (panel C). For the temperature AIS estimates, we show results for different values of $N_{\beta}$ in distinct colors. In panel C, we compare the online estimation in black with the offline Tr-AIS estimation, using the same models analyzed in panels A and B. Additionally, we present the error in LL estimation using a few machines and the trajectory parallel tempering (PTT), which will be detailed in the next section. Overall, the online Tr-AIS estimation consistently outperforms the other methods, and only the PTT estimation yields similar results in offline analyses.
Figure 3: Comparison of the error in (LL) estimation (with respect to the exact value) using different methods: AIS (left), AIS with a reference distribution (middle), and Tr-AIS (ours, right). For Tr-AIS, we present three approaches to computing it: online during training, offline using the models saved during training, and with PTT. The RBM was pretrained and trained with a total of 20 hidden nodes (allowing for an exact computation of the LL) for 10,000 gradient steps. PTT uses a subset of saved models such that the acceptance rate between two consecutive models is 0.25.
5.2 Standard Gibbs sampling vs. Parallel Trajectory Tempering
@@ 5.1 Trajectory Annealing importance sampling
One major challenge with structured datasets is quantifying the model’s quality, since sampling the equilibrium measure of a well-trained model is often too time-consuming. This affects the reliability of generated samples and indirect measures as log-likelihood’s estimation through Annealing Importance Sampling (AIS) (Krause et al., 2020), making them inaccurate and meaningless.
Figure 4: Comparison of the performance of different MCMC sampling methods on RBMs trained with the pretraining+PCD procedure on the MNIST01 (row-1), the HGD (row-2) and Ising 2D datasets (row-3). In A and B columns, we show the trajectory of two independent Markov chains (red and orange) iterated for $10^{4}$ MCMC steps using either PTT or alternate Gibbs sampling (AGS), projected onto the first two principal components of the dataset. The position of the chains is plotted every 10 MCMC steps. The black contour represents the density profile of the dataset. In column C, we show the average number of jumps between the two regions separated by the dashed grew line in the PCA plots using different MCMC methods: Alternative Gibbs Sampling (AGS), Parallel Tempering (PT) (Hukushima & Nemoto, 1996), without or with reference configuration (PTref) (Krause et al., 2020), the Stacked Tempering (Roussel et al., 2023), and the Trajectory Parallel Tempering (PTT) proposed in this work. The average is calculated over a population of 1000 chains and the shadow around the lines indicates the error of the mean.
@@
To illustrate this problem, let us consider the MNIST01, the HGD and the Ising 2D datasets. MNIST01 and Ising 2D are strongly bimodal in the PCA and the HGD trimodal, as shown in Fig. 1. Let us consider that we want to sample the equilibrium measure of the RBMs trained using low-rank RBM pretraining. In order to draw new samples from these models, one would typically run MCMC processes from random initialization and iterate them until convergence. The mixing time in such case is controlled by the jumping time between clusters. To accurately estimate the relative weight between modes, the MCMC processes must be ergodic, requiring many back-and-forth jumps. However, as shown in Figs. 4–A1-3, Alternate Gibbs sampling (AGS) dynamics are extremely slow, rarely producing jumps even after $10^{4}$ MCMC steps. The red curves in Figs. 4–C1-3 show the mean number of jumps over 1000 independent chains as a function of MCMC steps, indicating that a proper equilibrium generation would require at least $10^{6}-10^{7}$ MCMC steps. In order to speed up sampling, we consider other traditional optimized MCMC methods such as the PT (with or without reference), the Stack Tempering (Roussel et al., 2023) and our Parallel Trajectory Tempering (PTT) that will be explained below. The evolution of the number of jumps between modes observed with each of these algorithms, as a function of the total number of AGS steps performed (i.e. the number of parallel models multiplied by the number of AGS steps performed at each model) is shown in Fig. 4–C1-3. Before discussing our PTT algorithm and the results, let us briefly introduce the PT algorithm.
5.2.1 Parallel Tempering approaches
Parallel tempering (PT) is commonly used to simulate multiple systems in parallel at different inverse temperatures $\beta$ (Hukushima & Nemoto, 1996; Marinari & Parisi, 1992), with occasional temperature swaps between models based on the Metropolis rule. Since this move satisfies detailed balance, the dynamics will converge to the Boltzmann distribution $p_{\beta}(\bm{x})\propto\exp(-\beta\mathcal{H}(\bm{x}))$ at each $\beta$. However, a more general approach involves simulating multiple systems, each with its own set of parameters. We can define a system as $p_{\bm{\Theta}^{t}}(\bm{x})$, where $\bm{\Theta}^{t}$ denotes the parameters of the model indexed by $t$. This includes the classic PT case, where temperature scaling is applied, $\bm{\Theta}^{t}=\beta^{(t)}\bm{\Theta}$, with $\bm{\Theta}$ being the parameters of the model to be sampled.
In this work, we introduce a new algorithm that exchanges parameters of models saved at different stages of learning, using the learning trajectory to define the parameter set $\bm{\Theta}^{t}$, where $t$ refers to a specific point in the training process. We call this method Parallel Trajectory Tempering (PTT). The rule for swapping configurations $\bm{x}=(\bm{v},\bm{h})$ between neighboring machines indexed by $t$ and $t-1$ is given by
$p_{\mathrm{acc}}(\bm{x}^{t}\leftrightarrow\bm{x}^{t-1})=\min\left(1,\exp\left(\Delta\mathcal{H}^{t}(\bm{x}^{t})-\Delta\mathcal{H}^{t}(\bm{x}^{t-1})\right)\right),$ (5)
where again the index $t$ indicates that the Hamiltonian is evaluated using the parameters $\bm{\Theta}^{t}$. This move satisfies detailed balance with our target equilibrium distribution $p(\bm{x})=\exp(-\mathcal{H}(\bm{x}))/Z$, ensuring that the moves lead to the same equilibrium measure. The reason for using the trajectory is that PT algorithms work very well when a system undergoes a continuous phase transition. As discussed in the appendix G, adjusting the inverse temperature can lead to discontinuous transitions for which PT is not known to perform very well, while following the trajectory, the system undergoes several continuous transitions as discussed in Decelle et al. (2018); Bachtis et al. (2024).
Our PTT approach relies on the ease of sampling configurations from an initial RBM. In classical PCD training, the early “non-specialized” models thermalize quickly because they behave like Gaussian distributions. For RCM+PCD, independent configurations can be efficiently sampled from the low-rank RBM using a static Monte Carlo algorithm, as described in Section A.1 of the SI. The trajectory flow then significantly speeds up convergence to equilibrium, especially in the presence of a multimodal structure. The time interval between successive models is chosen to keep the acceptance probability for exchanges between neighboring models (Eq. 5) around 0.25 (we observed that the total time (number of AGS multiplied by number of machines) needed to thermalize a PTT chain did not depend on the acceptance fixed in the interval [0.2,0.5]. In practice, effective models can be saved during training by monitoring the acceptance rate between the last saved model and the current one. Pre-trained models require far fewer saved models, as many are naturally positioned near major phase transitions. The number of models used for each sampling process, along with a detailed description of the algorithm, is provided in the SI.
In Fig. 4–C1-3, we present the number of jumps between clusters as a function of the total number of AGS steps, where in PT-like schemes this refers to 1 (AGS step + one swap proposal) multiplied by the total number of models run in parallel. In all cases, PTT sampling outperforms the other methods, showing particularly remarkable results for HGD. Additionally, our method has the advantage of not requiring extra training like Stacked Tempering or simulating unnecessary temperatures.
Moreover, the equilibrium samples obtained from different models during sampling can be used to compute the log-likelihood, similar to the standard PT approach Krause et al. (2020). In Fig. 3–right we show that the quality of this estimation is comparable with the online trajectory AIS. Another desirable advantage of the PT algorithms is that one can easily check the thermalisation of the MCMC process by investigating the ergodicity of the random walk in models Banos et al. (2010). In addition, we evaluated other indicators of sample quality and potential overfitting. As detailed analysis is in Sect. H of the SI. Our method consistently provides high-quality samples across all datasets and, thanks to precise likelihood computation, can effectively detect overfitting.